# Planck’s Law from a Classical Free Energy Extremum Involving Fisher Information

### v2

@@ Abstract
We derive Planck’s law from a classical variational principle over probability densities, without invoking quantum states, quantized oscillator energies, or ensemble averages. We construct a generalized free energy functional involving entropy and Fisher information, with weights determined by the dimensionless ratio $\gamma=\hbar\omega/k_{B}T$. When extremized under a Gaussian ansatz, this functional yields the exact Planck distribution. The only quantum input is a minimal threshold assumption: that an oscillator emits a photon of energy $\hbar\omega$ only when a thermal fluctuation delivers at least that much energy. We also present a complementary kinetic derivation, based on threshold-activated thermal cascades, that yields the same result through classical stochastic reasoning. Together, these approaches suggest that Planck’s law — long considered a hallmark of quantum theory — may instead arise from classical thermodynamic principles supplemented by minimal constraints. This reframing has potential implications for understanding the emergence of quantum behavior from classical statistical systems.
@@ 1 Introduction
Our framework does not rely on quantum states, quantized oscillator levels, or Bose–Einstein statistics. The only explicitly quantum ingredient is the threshold assumption that photon emission requires energy $\hbar\omega$. The inclusion of Fisher information is motivated by information geometry and steady-state reasoning, and while it shares mathematical structure with quantum variational principles, it is employed here purely in a classical setting. This suggests an alternative physical foundation for black-body radiation, rooted in information geometry and thermodynamic variational principles. This direction complements earlier efforts to recover Planck’s law from classical frameworks, including models based on stochastic electrodynamics [5, 6] and statistical reinterpretations of Maxwell theory [7].
@@ 2.1 Entropy Scaling with Energy Quanta
$p=\mathbb{P}(E>\hbar\omega)=e^{-\gamma}.$
Since only a fraction $p$ of kicks are effective in producing emission, we argue that the entropy term in the free energy should be weighted by the relevant energy scale for emission events: $\hbar\omega$. Thus we set $f(\gamma)=\gamma$ since that makes the entropy term in the free energy $\hbar\omega H(p)$ rather than the typical $k_{B}TH(p).$ This ensures that entropy cost reflects the size of the required excitation, and is independent of the thermal environment’s effective temperature.
@@ 2.2 Fisher Information Scaling and Failed Kicks
Given that successful excitation occurs with probability $p=e^{-\gamma}$, the expected number of failed kicks per successful one is:
$n_{T}=\frac{1}{p}-1=e^{\gamma}-1.$
We propose that the number of failed kicks $n_{T}$ is proportional to the scale of the perturbation to $\nabla p(x)$, and hence $J(p)\propto n_{T}^{2}$. To preserve the interpretation of the Fisher term as a geometric constraint rather than a thermodynamic driver, we require that it not scale with temperature in the high-temperature limit. To prevent the Fisher term $g(\gamma)J[p]$ from scaling with temperature, we must then set $g(\gamma)\propto 1/n_{T}^{2}$. This condition determines the form of $g(\gamma)$ in our model:
$g(\gamma)+1\propto\frac{1}{n_{T}^{2}}=\frac{1}{(e^{\gamma}-1)^{2}}.$
The additive constant +1+1 ensures that in the zero-temperature limit, the Fisher coefficient tends to unity, so the Fisher term approaches $\hbar^{2}/8m\cdot J[p]$, as in the quantum free energy functional. However, the full functional only recovers the quantum variational principle when both $T\to 0$ and $\omega\to 0$ such that $\gamma=\hbar\omega/k_{B}T$ remains constant — in this joint limit, the entropy term vanishes and the quantum structure emerges exactly.
@@
The prefactor of 2 in the Fisher term ensures that in the high-temperature limit $(\gamma\ll 1)$, the variational free energy is minimized by a distribution whose variance reproduces the classical equipartition result: $\langle U\rangle=\frac{1}{2}k_{B}T$. Interestingly, even though our functional does not reduce to the quantum variational principle when $T\to 0$ at fixed $\omega$, the minimizing distribution still converges to the quantum mechanical ground state in this limit. This can be seen directly from Eqs. 8 and 9 below, which recover the quantum variance and zero-point energy as $T\to 0$, despite the presence of the entropy term in our free energy. Thus, while the functionals differ, their minima coincide in this limit. The function $g(\gamma)$ changes sign at $\gamma=\log 3$, corresponding to a crossover temperature $T_{c}=\hbar\omega/(k_{B}\log 3)$. For $\gamma<\log 3$, the Fisher term is positive and penalizes localization, consistent with thermal smoothing; for $\gamma>\log 3$, it becomes negative andfavors sharper localization, corresponding to reduced thermal fluctuations. This structure enables the Fisher term to encode the transition from thermal to quantum-dominated regimes. This assumes the oscillator samples energy fluctuations frequently enough to reach steady state, with strongest coupling near classical turning points. While no explicit dynamics are modeled, the steady-state variational principle may be viewed as characterizing the long-time average distribution resulting from thermal interactions.
@@
$\sigma^{2}=\frac{1}{2}\frac{\hbar}{m\omega}\left(1\pm\sqrt{1+g(\gamma)}\right)=\frac{1}{2}\frac{\hbar}{m\omega}\left(1+\frac{2}{e^{\gamma}-1}\right).$ (8)
Substituting into the expected potential energy yields
@@ Discussion
Finally, our kinetic interpretation—developed in Appendix A—provides an alternative classical route to the Planck’s law, based on thermal activation cascades. This framework suggests potential experimental signatures in low-temperature, single-mode thermal radiation systems. If energy enters electromagnetic modes in discrete bursts triggered by supra-threshold fluctuations, then photon emission may exhibit bunched statistics or non-Poissonian correlations. Future time-resolved measurements could test these predictions and clarify the microscopic dynamics underlying thermal radiation. A summary of major derivations, including our variational and kinetic formulations, is provided in Table 1. Our results suggest that Planck’s law may be viewed not as a direct consequence of quantization, but as the natural equilibrium outcome of classical systems governed by threshold constraints and variational structure. In particular, the cascade mechanism introduces the possibility of experimental signatures—such as photon clustering or correlated emission statistics—that could provide empirical tests of this classical activation-based model of radiation. More broadly, these results hint at a deeper connection between thermodynamic structure, information geometry, and quantum thermodynamics.
@@
This Work (Kinetic) Thermal kicks with $E>\hbar\omega$ trigger cascades Classical stochastic kinetics, detailed balance Threshold energy for emission only
@@ Acknowledgments
In this appendix, we present a classical kinetic model of radiation that offers an alternative derivation of Planck’s law. This approach is physically motivated by threshold-triggered energy cascades initiated by thermal fluctuations. We consider energy exchange between a thermal bath and a single electromagnetic (EM) mode of frequency $\omega$. The key physical mechanism is a stream of thermal interaction events, or “kicks,” delivered by atoms in the surrounding body. These kicks occur at an average rate $r$ that may depend on $T$, and their energies are drawn from a Boltzmann distribution with mean $k_{B}T$. Each kick represents a discrete opportunity for energy transfer, either into or out of the EM mode.
To inject energy into the mode, a kick must deliver at least $\hbar\omega$. The probability of this occurring is $p=e^{-\gamma}$, where $\gamma=\hbar\omega/k_{B}T$. A kick that exceeds this threshold initiates an emission cascade: a sequence of emissions continues as long as subsequent kicks also exceed the threshold, and ends with the first sub-threshold kick. This process produces a geometrically distributed number of emissions per cascade with mean $\langle N\rangle=1/(1-p)$, and average energy input per cascade $\langle E_{\text{cascade}}\rangle=\hbar\omega/(1-e^{-\gamma})$. Only a fraction $e^{-\gamma}$ of all kicks initiate cascades, so the average cascade rate is $\alpha=re^{-\gamma}$. The total energy inflow rate is then $\alpha\cdot\langle E_{\text{cascade}}\rangle=re^{-\gamma}\cdot\hbar\omega/(1-e^{-\gamma})$.
@@
A crucial aspect of this derivation is the presence of cascades. Without them, modeling energy input as a Poisson process where each supra-threshold kick contributes only one $\hbar\omega$ would yield $\langle E\rangle=\hbar\omega\cdot e^{-\gamma}$, which underestimates the true result. The cascades reflect the physical idea that a single large kick can unlock a temporary period of heightened coupling, during which multiple emissions occur in quick succession. This clustering of energy input events is essential to reaching the correct equilibrium distribution. Thus, Planck’s law emerges not from quantized field assumptions, but from classical thermodynamic reasoning involving threshold activation, stochastic fluctuation statistics, and asymmetry between input and output paths rooted in thermal accessibility. The kinetic picture complements the variational framework developed in the main text, and together they suggest a robust classical origin for the blackbody spectrum.
@@
Rewrite Eq. 4 in terms of the real-valued probability amplitude $\psi(x)$ via $p(x)=\psi(x)^{2}$. To extremize $F_{\gamma}[\psi]$, we introduce a Lagrange multiplier $\lambda$ to enforce normalization. The Lagrangian density $\mathcal{L}(\psi,\nabla\psi)$ becomes:
$U(x)\psi^{2}-k_{B}Tf(\gamma)\psi^{2}\log\psi^{2}+g(\gamma)\cdot\frac{\hbar^{2}}{2m}|\nabla\psi|^{2}-\lambda\psi^{2}.$
Applying the Euler–Lagrange equation, we obtain:
@@
$\left[-g(\gamma)\cdot\frac{\hbar^{2}}{2m}\nabla^{2}+U(x)+k_{B}Tf(\gamma)(1+2\log\psi(x))\right]\psi(x)$ (11)
$=\lambda\psi(x).$
This is the generalized Schrödinger equation arising from extremizing the free energy, with eigenvalue $\lambda$. The term with $\log\psi(x)$ introduces entropy-weighted nonlinearity, and $\lambda$ is the eigenvalue associated with the extremum. Setting $f(\gamma)=0$ and $g(\gamma)=1$ recovers Shrödinger, and choosing $f(\gamma)=1$ and $g(\gamma)=0$ recovers Boltzmann.
To compute explicit results for energy and variance, we now assume that the extremal solution $p(x)=\psi^{2}(x)$ is Gaussian with zero mean and variance $\sigma^{2}$. The free energy becomes:
$F(\sigma^{2})=\frac{1}{2}m\omega^{2}\sigma^{2}-\frac{1}{2}k_{B}Tf(\gamma)\log(2\pi e\sigma^{2})+\frac{g(\gamma)\hbar^{2}}{8m\sigma^{2}}.$
To find the extremum, we differentiate with respect to $\sigma^{2}$ and solve the resulting quadratic equation to obtain:
$\sigma^{2}=f(\gamma)\frac{1}{2}\frac{k_{B}T}{m\omega^{2}}\left(1+\sqrt{1+\frac{g(\gamma)\gamma^{2}}{f(\gamma)^{2}}}\right).$
Substitution of this Gaussian solution into Eq. 11 yields a free energy eigenvalue
@@
$\lambda$ $=$ $g(\gamma)\frac{\hbar^{2}}{4m\sigma^{2}}+f(\gamma)k_{B}T\left(1-\frac{1}{2}\log\left(2\pi\sigma^{2}\right)\right),$
which does not correspond to the physical energy. The eigenvalue $\lambda$ in Eq. 11 arises as a Lagrange multiplier enforcing the normalization constraint. While Eq. 11 resembles a generalized Schrödinger equation, its eigenvalue lacks a direct thermodynamic interpretation. Instead, the physically meaningful energy is computed from the extremal distribution $p(x)=\psi^{2}(x)$, using expectation values of the potential and (by equipartition) kinetic energy. In particular, the total oscillator energy $\langle E\rangle=\langle U\rangle+\langle T\rangle$ is obtained from the variance of the optimal Gaussian. This distinction is analogous to classical thermodynamics, where the minimum of a free energy functional determines equilibrium properties, but does not itself represent the internal energy.
For completeness, we confirm that the Gaussian ansatz indeed satisfies Eq. 11 by substituting $\psi(x)=\left(2\pi\sigma^{2}\right)^{-1/4}\exp\left(-\frac{x^{2}}{4\sigma^{2}}\right)$ into the left-hand side of Eq. 11, assuming a harmonic potential $U(x)=\frac{1}{2}m\omega^{2}x^{2}$. This gives
$\biggr{[}g(\gamma)\frac{\hbar^{2}}{2m}\left(\frac{1}{2\sigma^{2}}-\frac{x^{2}}{4\sigma^{4}}\right)+\frac{1}{2}m\omega^{2}x^{2}$
@@
$+k_{B}Tf(\gamma)\left(1-\frac{x^{2}}{2\sigma^{2}}-\frac{1}{2}\log(2\pi\sigma^{2})\right)\biggr{]}\psi(x)=\lambda\psi(x),$

### v3

@@ Abstract
We derive Planck’s law from a classical variational principle over probability densities, without invoking quantum states, quantized oscillator energies, or a canonical ensemble over discrete oscillator energy levels. We construct a generalized free energy functional involving entropy and Fisher information, with weights determined by the dimensionless ratio $\gamma=\hbar\omega/k_{B}T$. When extremized under a Gaussian ansatz, this functional yields the exact Planck distribution. The only quantum input is a minimal threshold assumption: that an oscillator emits a photon of energy $\hbar\omega$ only when a thermal fluctuation delivers at least that much energy. We also present a complementary kinetic derivation, based on threshold-activated thermal emission cascades, that yields the same result through classical stochastic reasoning. Together, these approaches suggest that Planck’s law — long considered a hallmark of quantum theory — may instead arise from classical thermodynamic principles supplemented by minimal constraints. This reframing has potential implications for understanding the emergence of quantum behavior from classical statistical systems.
@@ 1 Introduction
Our framework does not rely on quantum states, quantized oscillator levels, or Bose–Einstein statistics. The only explicitly quantum ingredient is the threshold assumption that photon emission requires energy $\hbar\omega$. The present framework does not attempt to derive the existence of photons or the relation $E_{\gamma}=\hbar\omega$ from purely classical dynamics. Instead, we treat $E_{\gamma}=\hbar\omega$ as an experimentally established constraint on radiative emission and ask a narrower question: given this minimal threshold, can the full Planck factor $(e^{\gamma}-1)^{-1}$ emerge from a classical free-energy extremum over continuous densities without quantized oscillator levels or Bose counting? Our inclusion of Fisher information is motivated by information geometry and steady-state reasoning, and while it shares mathematical structure with quantum variational principles, it is employed here purely in a classical setting. This suggests an alternative physical foundation for black-body radiation, rooted in information geometry and thermodynamic variational principles. [^1] We use information geometry to mean the standard view of probability densities as a statistical manifold whose local metric is induced by the Fisher information, which measures localization/roughness of probability densities in configuration space. We refer to the Fisher term in Eq. (4) as a geometric constraint because it penalizes the shape of the steady-state density rather than representing a thermal driving contribution. This direction complements earlier efforts to recover Planck’s law from classical frameworks, including models based on stochastic electrodynamics [2, 8] and statistical reinterpretations of Maxwell theory [9].
@@ 2.1 Entropy Scaling with Energy Quanta
$p_{\mathrm{th}}=\mathbb{P}(E>\hbar\omega)=e^{-\gamma}.$
Since only a fraction $p_{\mathrm{th}}$ of kicks are effective in producing emission, we argue that the entropy term in the free energy should be weighted by the relevant energy scale for emission events: $\hbar\omega$. Thus we set $f(\gamma)=\gamma$ since that makes the entropy term in the free energy $\hbar\omega H(p)$ rather than the typical $k_{B}TH(p).$ This ensures that entropy cost reflects the size of the required excitation, and is independent of the thermal environment’s effective temperature.
@@ 2.2 Fisher Information Scaling and Failed Kicks
Given that successful excitation occurs with probability $p_{\mathrm{th}}=e^{-\gamma}$, the expected number of failed kicks per successful one is:
$n_{T}=\frac{1}{p_{\mathrm{th}}}-1=e^{\gamma}-1.$
We propose that the number of failed kicks $n_{T}$ is proportional to the scale of the perturbation to $\nabla p(x)$, and hence $J(p)\propto n_{T}^{2}$. To preserve the interpretation of the Fisher term as a geometric constraint rather than a thermodynamic driver, we require that it not scale with temperature in the high-temperature limit. To avoid an unphysical growth of the Fisher contribution in the high-temperature limit, we take the Fisher prefactor to scale inversely with the square of the expected number of sub-threshold (failed) attempts per successful threshold crossing. Since $n_{T}=e^{\gamma}-1$, this motivates a scaling $1+g(\gamma)\propto(e^{\gamma}-1)^{-2}$. We choose the constant offset so that $g(\gamma)\geq-1$ and the extremal solution approaches the $T\to 0$ ground-state variance $\sigma^{2}\to\hbar/(2m\omega)$, while still reproducing the correct thermal scaling for $\gamma\ll 1$.
@@
The prefactor of 2 in the Fisher term ensures that in the high-temperature limit $(\gamma\ll 1)$, the variational free energy is minimized by a distribution whose variance reproduces the classical equipartition result $\langle U\rangle=\frac{1}{2}k_{B}T$, up to the $T$-independent zero-point offset. Interestingly, even though our functional does not reduce to the quantum variational principle when $T\to 0$ at fixed $\omega$, the minimizing distribution still converges to the quantum mechanical ground state in this limit. This can be seen directly from Eqs. 8 and 9 below, which recover the quantum variance and zero-point energy as $T\to 0$, despite the presence of the entropy term in our free energy. Thus, while the functionals differ, their minima coincide in this limit. The function $g(\gamma)$ changes sign at $\gamma=\log 3$, corresponding to a crossover temperature $T_{c}=\hbar\omega/(k_{B}\log 3)$. For $\gamma<\log 3$, the Fisher term is positive and penalizes localization, consistent with thermal smoothing; for $\gamma>\log 3$, it becomes negative and favors sharper localization, corresponding to reduced thermal fluctuations. This structure enables the Fisher term to encode the transition from thermal to quantum-dominated regimes. This assumes the oscillator samples energy fluctuations frequently enough to reach steady state, with strongest coupling near classical turning points. While no explicit dynamics are modeled, the steady-state variational principle may be viewed as characterizing the long-time average distribution resulting from thermal interactions.
@@
$\sigma^{2}=\frac{1}{2}\frac{\hbar}{m\omega}\left(1\pm\sqrt{1+g(\gamma)}\right)=\frac{1}{2}\frac{\hbar}{m\omega}\left(1+\frac{2}{e^{\gamma}-1}\right),$ (8)
where we take the positive branch to ensure that $\sigma^{2}>0,$ as well as the appropriate high temperature limit. Substituting into the expected potential energy yields
@@ Discussion
This Work (Kinetic) Thermal kicks with $E>\hbar\omega$ trigger emission cascades Classical stochastic kinetics, detailed balance Threshold energy for emission only
@@
Finally, our kinetic interpretation—developed in Appendix A—provides an alternative classical route to the Planck’s law, based on thermal emission cascades. An emission cascade is a burst process in which an above-threshold activation event opens a short-lived “emissive” window. During this window, emission attempts occur repeatedly and each attempt succeeds only if the instantaneous thermal energy transfer exceeds $\hbar\omega$; the window closes at the first sub-threshold attempt. This produces a geometrically distributed burst size (Appendix D provides a minimal stochastic realization and simulation). This framework suggests potential experimental signatures in low-temperature, single-mode thermal radiation systems. If energy enters electromagnetic modes in discrete bursts triggered by supra-threshold fluctuations, then photon emission may exhibit bunched statistics or non-Poissonian correlations. Future time-resolved measurements could test these predictions and clarify the microscopic dynamics underlying thermal radiation. A summary of major derivations, including our variational and kinetic formulations, is provided in Table 1. Our results suggest that Planck’s law may be viewed not as a direct consequence of quantization, but as the natural equilibrium outcome of classical systems governed by threshold constraints and variational structure. In particular, the cascade mechanism introduces the possibility of experimental signatures—such as photon clustering or correlated emission statistics—that could provide empirical tests of this classical activation-based model of radiation. More broadly, these results hint at a deeper connection between thermodynamic structure, information geometry, and quantum thermodynamics.
@@ Acknowledgments
In this appendix, we present a classical kinetic model of radiation that offers an alternative derivation of Planck’s law. This approach is physically motivated by threshold-triggered emission cascades initiated by thermal fluctuations. We consider energy exchange between a thermal bath and a single electromagnetic (EM) mode of frequency $\omega$. The key physical mechanism is a stream of thermal interaction events, or “kicks,” delivered by atoms in the surrounding body. These kicks occur at an average rate $r$ that may depend on $T$, and their energies are drawn from a Boltzmann distribution with mean $k_{B}T$. Each kick represents a discrete opportunity for energy transfer, either into or out of the EM mode. The kinetic balance model here describes the thermally activated, radiative energy stored in the mode, i.e. the temperature-dependent part above the $T\to 0$ baseline.
To inject energy into the EM mode, a kick must deliver at least $\hbar\omega$. The probability of this occurring is $p_{\mathrm{th}}=e^{-\gamma}$, where $\gamma=\hbar\omega/k_{B}T$. A kick that exceeds this threshold initiates an emission cascade: a sequence of emissions continues as long as subsequent kicks also exceed the threshold, and ends with the first sub-threshold kick. Equivalently, we idealize an above-threshold kick as opening a short-lived emissive window during which multiple emission attempts occur; each attempt independently exceeds threshold with probability $e^{-\gamma}$, and the window closes on the first sub-threshold attempt. This process produces a geometrically distributed number of emissions per cascade with mean $\langle N\rangle=1/(1-p_{\mathrm{th}})$, and average energy input per cascade $\langle E_{\text{cascade}}\rangle=\hbar\omega/(1-e^{-\gamma})$. Only a fraction $e^{-\gamma}$ of all kicks initiate cascades, so the average cascade rate is $\alpha=re^{-\gamma}$. The total energy inflow rate is then $\alpha\cdot\langle E_{\text{cascade}}\rangle=re^{-\gamma}\cdot\hbar\omega/(1-e^{-\gamma})$.
@@
A crucial aspect of this derivation is the presence of emission cascades.Without cascades, suppose each kick injects at most one quantum: with probability $p_{\mathrm{th}}=e^{-\gamma}$ a kick adds energy $\hbar\omega$, otherwise it adds zero. Then the mean input power is $r\,p_{\mathrm{th}}\,\hbar\omega$, while under the same linear-loss assumption the mean absorption power is $r\,\langle E\rangle$. Steady-state balance gives $\langle E\rangle=\hbar\omega\,e^{-\gamma},$ which underestimates the Planck mean; the geometric cascades are precisely what supply the missing multiplicative factor $(1-e^{-\gamma})^{-1}$. The cascades reflect the physical idea that a single large kick can unlock a temporary period of heightened coupling, during which multiple emissions occur in quick succession. This clustering of energy input events is essential to reaching the correct equilibrium distribution. Thus, Planck’s law emerges not from quantized field assumptions, but from classical thermodynamic reasoning involving threshold activation, stochastic fluctuation statistics, and asymmetry between input and output paths rooted in thermal accessibility. The kinetic picture complements the variational framework developed in the main text, and together they suggest a robust classical origin for the blackbody spectrum.
@@
We start from our free energy functional $F_{\gamma}[p]$ in Eq. (4). To streamline the Euler–Lagrange computation, we rewrite the functional in terms of a real probability amplitude $\psi(x)\geq 0$ defined by
$p(x)=\psi(x)^{2},\qquad\int\psi(x)^{2}\,dx=1.$
In one dimension, the Fisher information $J[p]$ becomes
$J[p]=\int\frac{(2\psi\,\partial_{x}\psi)^{2}}{\psi^{2}}\,dx=4\int(\partial_{x}\psi)^{2}\,dx,$
so that
$\frac{\hbar^{2}}{8m}J[p]=\frac{\hbar^{2}}{2m}\int(\partial_{x}\psi)^{2}\,dx.$
Also, since $H[p]=\int p\log p\,dx$, the entropy contribution becomes
$k_{B}Tf(\gamma)H[p]=k_{B}Tf(\gamma)\int\psi(x)^{2}\log\!\bigl(\psi(x)^{2}\bigr)\,dx.$
Therefore, the functional can be written as
@@
$F_{\gamma}[\psi]=\int\!\left[U(x)\psi(x)^{2}+k_{B}Tf(\gamma)\,\psi(x)^{2}\log\!\bigl(\psi(x)^{2}\bigr)+g(\gamma)\,\frac{\hbar^{2}}{2m}\,(\partial_{x}\psi(x))^{2}\right]dx.$ (11)
To keep $p(x)$ normalized, we impose normalization with a Lagrange multiplier $\lambda$:
$\widetilde{F}_{\gamma}[\psi]=F_{\gamma}[\psi]-\lambda\left(\int\psi(x)^{2}\,dx-1\right).$
Next, we apply the Euler–Lagrange (EL) equation $\frac{\partial\mathcal{L}}{\partial\psi}-\partial_{x}\!\left(\frac{\partial\mathcal{L}}{\partial(\partial_{x}\psi)}\right)=0$ to the integrand of $\widetilde{F}_{\gamma}[\psi],$ which ignoring the constant $\lambda$ is:
$\mathcal{L}(\psi,\partial_{x}\psi)=U\psi^{2}+k_{B}Tf(\gamma)\,\psi^{2}\log(\psi^{2})+g(\gamma)\,\frac{\hbar^{2}}{2m}\,(\partial_{x}\psi)^{2}-\lambda\psi^{2}.$
Since
$\frac{\partial}{\partial\psi}\bigl(\psi^{2}\log(\psi^{2})\bigr)=2\psi\bigl(1+\log(\psi^{2})\bigr)=2\psi\bigl(1+2\log\psi\bigr),$
the first derivative in the EL equation is
$\frac{\partial\mathcal{L}}{\partial\psi}=2U\psi+2k_{B}Tf(\gamma)\,\psi\bigl(1+2\log\psi\bigr)-2\lambda\psi.$
Similarly,
@@
$\frac{\partial\mathcal{L}}{\partial(\partial_{x}\psi)}=g(\gamma)\,\frac{\hbar^{2}}{m}\,\partial_{x}\psi,\qquad\partial_{x}\!\left(\frac{\partial\mathcal{L}}{\partial(\partial_{x}\psi)}\right)=g(\gamma)\,\frac{\hbar^{2}}{m}\,\partial_{x}^{2}\psi.$
Substituting and dividing by $2$ yields the generalized stationary equation:
@@
$g(\gamma)\,\frac{\hbar^{2}}{2m}\,\partial_{x}^{2}\psi(x)+\left[U(x)+k_{B}Tf(\gamma)\,\bigl(1+2\log\psi(x)\bigr)\right]\psi(x)=\lambda\,\psi(x).$ (12)
Equation (12) plays the role of a generalized stationary Schrödinger equation. Here $\lambda$ is the Lagrange multiplier enforcing normalization, and therefore should be interpreted as a free-energy eigenvalue rather than as the physical oscillator energy. This distinction matters because the entropy term introduces a nonlinear dependence on $\psi$, so $\lambda$ does not coincide with $\langle E\rangle$ even in the harmonic case. In our framework, physically meaningful energies are instead computed directly from expectation values under the extremal distribution (as in the main text), rather than inferred from $\lambda$.
To compute explicit results for the harmonic oscillator, we now assume the extremal solution $p(x)=\psi^{2}(x)$ is Gaussian with zero mean and variance $\sigma^{2}$. Using the standard Gaussian values for $H[p]$ and $J[p]$, and substituting into Eq. (4), the free energy reduces to
$F(\sigma^{2})=\frac{1}{2}m\omega^{2}\sigma^{2}-\frac{1}{2}k_{B}Tf(\gamma)\,\log(2\pi e\,\sigma^{2})+g(\gamma)\,\frac{\hbar^{2}}{8m}\,\frac{1}{\sigma^{2}}.$ (13)
To extremize this functional, we differentiate Eq. (13) with respect to $\sigma^{2}$, set the result to zero, and obtain a quadratic in $y=\sigma^{2}$, whose positive root yields
$\sigma^{2}=\frac{k_{B}Tf(\gamma)}{2m\omega^{2}}\left[1+\sqrt{1+\frac{g(\gamma)\hbar^{2}\omega^{2}}{k_{B}^{2}T^{2}f(\gamma)^{2}}}\right].$ (14)
We select the ++ branch because it reproduces the classical high-temperature scaling $\sigma^{2}\sim k_{B}T/(m\omega^{2})$.
Next, we specialize Eq. (14) to $f(\gamma)=\gamma$ and $\gamma=\hbar\omega/(k_{B}T)$, which implies
$\sigma^{2}=\frac{\hbar}{2m\omega}\left[1+\sqrt{1+g(\gamma)}\right].$
Using Eq. (5) (so that $1+g(\gamma)=\bigl(\tfrac{2}{e^{\gamma}-1}\bigr)^{2}$) gives Eq. (8). The mean potential energy in Eq. (9) follows by direct substitution of $\sigma^{2}$ from Eq. (8) into $\langle U\rangle=\frac{1}{2}m\omega^{2}\sigma^{2}$. Finally, invoking the symmetric phase-space extension described in the main text (so that $\langle T\rangle=\langle U\rangle$), the total mean oscillator energy (Eq. (10)) is $\langle E\rangle=\frac{\hbar\omega}{2}+\frac{\hbar\omega}{e^{\gamma}-1}$, and subtracting the $T\to 0$ limit isolates the radiative portion $\langle E_{\mathrm{rad}}\rangle=\frac{\hbar\omega}{e^{\gamma}-1}$, which is exactly the Planck mean energy per mode used in Eqs. (1)–(2).
Consistency check and the eigenvalue $\lambda$.
For completeness, we confirm that the Gaussian ansatz is consistent with the generalized stationary equation Eq. (12). Substituting $\psi(x)=(2\pi\sigma^{2})^{-1/4}\exp\!\left(-\frac{x^{2}}{4\sigma^{2}}\right)$ and $U(x)=\frac{1}{2}m\omega^{2}x^{2}$ into the left-hand side of Eq. (12) gives
$\biggl[\,$ $g(\gamma)\,\frac{\hbar^{2}}{2m}\left(\frac{1}{2\sigma^{2}}-\frac{x^{2}}{4\sigma^{4}}\right)+\frac{1}{2}m\omega^{2}x^{2}$
$\quad+k_{B}Tf(\gamma)\left(1-\frac{x^{2}}{2\sigma^{2}}-\frac{1}{2}\log(2\pi\sigma^{2})\right)\biggr]\psi(x)=\lambda\,\psi(x).$
Collecting the $x^{2}$ terms inside the brackets yields the coefficient
$B=\frac{1}{2}m\omega^{2}-\frac{1}{2}k_{B}Tf(\gamma)\frac{1}{\sigma^{2}}-g(\gamma)\frac{\hbar^{2}}{8m}\frac{1}{\sigma^{4}}.$
Since $\lambda$ is $x$-independent, Eq. (12) holds for all $x$ if and only if $B=0$. But differentiating Eq. (13) with respect to $\sigma^{2}$ gives $\frac{dF}{d(\sigma^{2})}=B$, so this EL consistency condition is exactly the same as the extremality condition used to obtain Eq. (14). With $B=0$, the remaining constant term fixes the Lagrange multiplier:
$\lambda=g(\gamma)\frac{\hbar^{2}}{4m\sigma^{2}}+f(\gamma)k_{B}T\left(1-\frac{1}{2}\log\left(2\pi\sigma^{2}\right)\right).$